A mutual fund might produce 5% return, with 40% probability, and 4% return, with 60% probability. Thus the standard deviation of this asset is: " with steps"

 

a.1.67%

b.0.49%

c.0.34%

d.0.65%

e.0.23%


1
Expert's answer
2022-02-13T12:19:10-0500

"5\times0.4+4\times0.6=4.4"


"standard deviation=(\frac{(5-4.4)^2+(4-4.4)^2}{2-1})^{0.5}=0.51"